A body is thrown from the surface of the earth with a velocity of  u m/s. The maximum height in meters reached by it will be ? (Answer should be in terms of R ,u ,g where R is radius of earth and,g is acceleration due to gravity)

Grade:12

2 Answers

Arun
4 years ago
Dear student
 
Maximum height = u² sin²theta/2g = u² /2g
 
Niw maximum R =u² /g
 
Hence maximum height = R/2
Vikas TU
3 years ago
Dear student 
Above ans is incorrect 
Applying conservation of mechanical energy 
1/2 mu^2 - GMm/R = - GMm/(R+h)
Solving this 
R+h = 2gR^2 /(2gR -u^2)
h = 2gR^2 /(2gR-u^2) - R 
h = u^2/(gR-u^2)
Good Luck